yourgrandpa:

members of a band estimated that 200 people were at their show last night. their manager told them that 227 people had actually been at the show. (what was the band's percent error?)

Vocaloid:

Calculate (estimated # of people - actual # of people) / (actual # of people) * 100%

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid Calculate (estimated # of people - actual # of people) / (actual # of people) * 100% \(\color{#0cbb34}{\text{End of Quote}}\) 200?

Vocaloid:

200 is your estimated 227 is your actual (estimated # of people - actual # of people) / (actual # of people) * 100% Therefore becomes (200 - 227) / (227) * 100%

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid 200 is your estimated 227 is your actual (estimated # of people - actual # of people) / (actual # of people) * 100% Therefore becomes (200 - 227) / (227) * 100% \(\color{#0cbb34}{\text{End of Quote}}\) 2700 ?

yourgrandpa:

@Vocaloid or is it 100?? kinda confused

Vocaloid:

0 200 - 227 is -27 -27 / 227 = -0.119 -0.119 * 100% = -11.9%

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid 0 200 - 227 is -27 -27 / 227 = -0.119 -0.119 * 100% = -11.9% \(\color{#0cbb34}{\text{End of Quote}}\) so its -11.9%

Vocaloid:

Yes

yourgrandpa:

ok thanks alot(: